OpenStudy (anonymous):

What is the equation of the line that passes through (-3, -1) and has a slope of 3/5? Put your answer in slope-intercept form. y = 3/5x + 4/5 y = 3/5x - 4/5 y = -3/5x - 4/5

OpenStudy (anonymous):

y=3/5x+4/5

OpenStudy (anonymous):

It should be A. It has a slope of 3/5, so the third one is right out. Solving for b in the equation y = ax + b: -1 = (3/5)(-3) + b -1 + 9/5 = b So b = 4/5

OpenStudy (anonymous):

And the equation becomes y = (3/5)x + 4/5